OpenStudy (anonymous):

given the functions f(x) = cos x and g(x) = 2^x+5. determine the value of g(f(pi))

OpenStudy (mertsj):

\[g(f(x))=2^{\cos x}+5\]

OpenStudy (mertsj):

\[g(f(\pi))=2^{\cos \pi}+5\]

OpenStudy (mertsj):

\[g(f(\pi))=2^{-1}+5\]

OpenStudy (mertsj):

\[g(f(\pi))=\frac{1}{2} +5=\frac{11}{2}\]